OpenStudy (anonymous):

simplify the expression 8-4s+6+8s=

OpenStudy (anonymous):

4s+14

OpenStudy (gw2011):

8+6=14 8s-4s=4s 4s+14=2(2s+7)

OpenStudy (anonymous):

First, separate the numbers and the variables, so you get 8 + 6 - 4s + 8s. Add the numbers together and the variables together, to get 14 + 4s. Now, you can factor 2 out because 14 and 4s have a greatest common factor of 2, to get 2 * (7 + 2s).